An investment of ​$106,000 was made by a business club. The investment was split into three parts and lasted for one year. The first part of the investment earned​ 8% interest, the second​ 6%, and the third​ 9%. Total interest from the investments was $8100. The interest from the first investment was 4 times the interest from the second. Find the amounts of the three parts of the investment.
Answer by mananth(16946)

An investment of ​$106,000 was made by a business club.
The investment was split into three parts and lasted for one year.
The first part of the investment earned​ 8% interest, Let investment be x
the second​ 6%, let investment be y
and the third​ 9%. let investment be z

x+y+z =106000----------------------------------------(1)
Total interest from the investments was $8100.
8%x+6%y+9%z = 8100 -------------------------------(2)

The interest from the first investment was 4 times the interest from the second.
8%x= 4*6%y
8x=24y
x=3y
substitute x=3y in (1)
x+y+z=106000
4y+z=106000 ----------------------------(3)
8%x+6%y+9%z = 8100
substitute x=3y
8%(3y)+6%y+9%z=8100
Multiply by 100
24y +6y +9z= 810000
30y+9z= 810000 --------------------------(4)
multiply (3)by 9 and subtract (4)
36y+9z=954000
30y+9z= 810000
6y= 144000
y= 24000
x=3y
x= 72000
106000 -96000 = 10000 =z
The first part of the investment earned​ 8% interest. $72,000
the second​ 6%, investment . $24,000
and the third​ 9%. investment $ 10,000
Check
72000*8%+24000*6%+9%*10000 =8100